What is the ph of stomach acid and the acid produced by stomach cells
1
Expert's answer
2020-12-08T05:29:45-0500

in different departments of the gastrointestinal tract is different:

The maximum theoretically possible acidity in the stomach: pH = 0.86 (corresponds to the acid production of 160 mmol / l).

The minimum theoretically possible acidity in the stomach: pH = 8.3 (corresponds to the pH of a saturated solution of HCO3-ions).

Normal acidity in the lumen of the stomach body on an empty stomach: pH = 1.5 — 2.0.

Acidity on the surface of the epithelial layer facing the lumen of the stomach: pH = 1.5-2.0.

Normal acidity in the antral part of the stomach: pH = 1.3-7.4.

Acidity in the depth of the epithelial layer of the stomach: about pH = 7.0.

Normal acidity in the esophagus: pH = 6.0-7.0.

Normal acidity in the duodenal bulb: pH = 5.6 — 7.9.

The acidity of the small intestine juice: 7.2-7.5 pH; with increased secretion, it reaches pH = 8.6.

The acidity of the large intestine: pH = 8,5 — 9,0.